OpenStudy (anonymous):

how to integrate : (cos(sx)/(1+x^2))dx

OpenStudy (anonymous):

\[\int\limits \frac{ \cos(sx) }{ 1+x^2 }ds\] you can treat 1/(1+x^2) as a constant

OpenStudy (anonymous):

Let u = sx, du = xds, can you go from there?

OpenStudy (anonymous):

\[\frac{ 1 }{ (1+x^2) } \int\limits \cos(sx) ds\]

OpenStudy (anonymous):

Hey did you figure out, still need help?

OpenStudy (anonymous):

iam sorry ... a slight mistake in the question .... it is not 'ds'.. but ' dx'